OpenStudy (anonymous):

Evaluate. 2^5/2^3 A)1 B)2^8 C)4 D) 1/4

OpenStudy (anonymous):

Okay, so 2 * 2 * 2 * 2 * 2 is what? 32 right? so you now have 32/2^3. 2*2*2 is what? 8 right? Now you have 32/8. 32/8 is 4. That is the answer. Make sense?

OpenStudy (anonymous):

Oh okay yeah i do! Thanks! :)

OpenStudy (anonymous):

2^5=2*2*2*2*2= 34 2^3=2*2*2= 8

OpenStudy (anonymous):

2 * 2 =4. 4 * 2 = 8. 8 * 2 = 16. 16 * 2 = 32. 2 * 2 = 4. 4 * 2 = 8. There ya go! You are welcome! Here is a medal if you can do this one. 2 ^ 3 ______ 2 ^ 2

OpenStudy (anonymous):

@Kyliejones

OpenStudy (anonymous):

an easy way is to subtract the exponents. so for his, subtract 2 from 3 and your left with 2^1 which equals what?

OpenStudy (anonymous):

Sorry i ran to get something im back ok and lets see the top one is 2^3 is = 8 2^2 is = 4

OpenStudy (anonymous):

So divide them and get.......@Kyliejones

OpenStudy (anonymous):

4 goes into 8 how many times?

OpenStudy (anonymous):

@Kyliejones

OpenStudy (anonymous):

2